OpenStudy (anonymous):

Solve |2x + 8| = 6. Show its equivalent compound sentence and show its solution. Solve |2x + 8| = 6. Show its equivalent compound sentence and show its solution. @Mathematics

OpenStudy (anonymous):

\[2x+8=6\text { or } 2x+8=-6\]

jhonyy9 (jhonyy9):

2x +8 = 6 2x= -2 x= -1 2x+8= -6 2x= -14 x= -7